OpenStudy (anonymous):

Does anyone know how to simplify this expression: (3x^-3/4x^3)^-2 with a bit of explanation?

OpenStudy (turingtest):

combine the following rules\[\frac{x^a}{x^b}=x^{a-b}\]\[(x^a)^b=x^{ab}\]